OpenStudy (cookie_2046):

evaluate the expression if a=1 and b=4

OpenStudy (cookie_2046):

evaluate the expression if a=1 and b=4 \[\frac{ 2(a+6b }{ ab}\] A. 12.5 B.8 C.14 D. 6.5

OpenStudy (solomonzelman):

First will make sure we know that: \(\large\color{slate}{ \displaystyle \frac{2(a+6b)}{ab} }\) is same as \(\large\color{slate}{ \displaystyle \frac{2(a+6\cdot b)}{a\cdot b} }\)

OpenStudy (solomonzelman):

now, when a=1 and b=4 you have: \(\large\color{slate}{ \displaystyle \frac{2(1+6\cdot 4)}{1\cdot 4} }\)

OpenStudy (solomonzelman):

just simplify this....

OpenStudy (cookie_2046):

ok

OpenStudy (solomonzelman):

when you are done tell me what you got pliz.

OpenStudy (cookie_2046):

The answer is (A) 12.5 Correct? @SolomonZelman

OpenStudy (solomonzelman):

yes, 12.5 is correct.

OpenStudy (cookie_2046):

Ok thanks for helping

OpenStudy (solomonzelman):

anytime